OpenStudy (anonymous):

The radius of a sphere is divided by 3. What is the effect on the volume? will v be divided by 3?

OpenStudy (anonymous):

v = [4 (pi)(r)^3] /3 If r is divided by 3, v = [4 (pi)(r/3)^3]/3 v = [4 (pi)((r)^3) / 9]/ 3 v = [4(pi)(r)^3] / 27 v = So the volume will be divided by 9 since when the original eqn is divided by the new one, you get that.